# Exploring Scene Coherence for Semi-Supervised 3D Semantic Segmentation

### v1

@@ Abstract
Semi-supervised semantic segmentation, which efficiently addresses the limitation of acquiring dense annotations, is essential for 3D scene understanding. Most methods leverage the teacher model to generate pseudo labels, and then guide the learning of the student model on unlabeled scenes. However, they focus only on points with pseudo labels while directly overlooking points without pseudo labels, namely intra-scene inconsistency, leading to semantic ambiguity. Moreover, inter-scene correlation between labeled and unlabeled scenes contribute to transferring rich annotation information, yet this has not been explored for the semi-supervised tasks. To address these two problems, we propose to explore scene coherence for semi-supervised 3D semantic segmentation, dubbed CoScene. Inspired by the unstructured and unordered nature of the point clouds, our CoScene adopts the straightforward point erasure strategy to ensure the intra-scene consistency. Moreover, patch-based data augmentation is proposed to enhance the inter-scene information transfer between labeled and unlabeled scenes at both scene and instance levels. Extensive experimental results on SemanticKITTI and nuScenes show that our approach outperforms existing methods.
Introduction
LiDAR (Light Detection and Ranging) point cloud semantic segmentation, aiming to recognize the class label for each point of an input point cloud scene, is a core task in autonomous driving. Due to the advancement of deep neural networks, tremendous progress has been achieved (Hu et al. 2020; Zhu et al. 2021; Su et al. 2023). However, the excellent performance of most studies depends heavily on large-scale datasets with high-quality point-level annotations. Such dense and delicate point-level labeling is labor-intensive and time-consuming, hindering the deployment of point cloud semantic segmentation models in real-world applications. Therefore, there is a compelling need to reduce the dependence on huge exhaustively labeled datasets.
Figure 1: (a) Generated pseudo-label scenes from unlabeled scenes containing points with and without pseudo labels simultaneously (green and gray). (b) Regardless of whether point-based or voxel-based segmentation networks are employed, resulting in inconsistencies between forward and backward propagation. (c) Annotation information transfer between labeled and unlabeled scenes.
Semi-supervised learning (SSL), a significant machine learning paradigm for alleviating data annotation demands, seeks to utilize a small number of labeled samples alongside a vast collection of unlabeled samples to optimize deep models. Recently, several semi-supervised 3D semantic segmentation methods (Kong et al. 2023b; Jiang et al. 2021; Li, Shum, and Breckon 2023) have been proposed, with a core focus on effectively transferring diverse semantic information. These methods typically first obtain pseudo-labeled scenes from unlabeled data using the teacher network, and then feed the entire pseudo-labeled scenes into the student network for training. Although these methods have reported promising results, two main issues still need to be addressed.
As shown in Figure 1 (a), the pseudo-labeled scene consists of points with and without pseudo-labels simultaneously. During training, all points are involved in the model’s forward propagation, while only points with pseudo-labels are used for backward propagation, and other points are overlooked. Inconsistency between forward and backward propagation within the scene leads to semantic ambiguity, meaning the segmentation model is unable to determine the semantic labels for points without pseudo-labels. Consequently, this limits the effectiveness of semantic information transfer. Both point-based (Hu et al. 2020) and voxel-based (Choy, Gwak, and Savarese 2019; Zhu et al. 2021) methods suffer from this issue, as illustrated in Figure 1 (b).
Diverse semantic information is crucial for the success of semi-supervised learning of the segmentation models (Kong et al. 2023b; Xiao et al. 2022). Thus, many studies focus on developing data augmentation methods (Sheshappanavar, Singh, and Kambhamettu 2021; Zhang et al. 2022) to expand the data distribution. However, these methods typically model the correlation between two scenes, while neglecting the correlation among multiple scenes. As shown in Figure 1 (c), the labeled and unlabeled scenes may exhibit similarities in factors such as spatial structures. Exploring the correlations between scenes can be beneficial for annotation information transfer, thereby enhancing learning on unlabeled scenes.
To address the above two issues, we explore Scene Coherence and propose CoScene for semi-supervised 3D semantic segmentation. Specifically, based on the unstructured and unordered nature of point clouds, we design a point erasure strategy that removes the points without pseudo labels during forward propagation, achieving full annotation of the training scene and ensuring intra-scene consistency. This strategy complements the insufficiency of the basic pseudo-label mechanism, which can serve as a plug-and-play module for semi-supervised 3D segmentation frameworks based on pseudo labels. Additionally, we design a patch-based data augmentation to model the correlation between multiple scenes at both the scene and instance levels. The key idea is to construct two patch pools to store scene-level and instance-level information, and then combine scene-level and instance-level patches from multiple scenes to generate diverse training samples, facilitating the transfer of annotation information to unlabeled scenes.
The proposed CoScene has been extensively evaluated through experiments and has demonstrated superiority across various benchmarks for semi-supervised 3D semantic segmentation. Under the setting of 5% labeled scenes, our method outperforms the state-of-the-art method by 4% in terms of mIoU on the SemanticKITTI (Behley et al. 2019) dataset. To summarize, our contributions are as follows:
@@ Introduction
  • We propose a plug-and-play point erasure strategy complementing with pseudo-label mechanism in semi-supervised point cloud semantic segmentation, ensuring intra-scene consistency.
  • We propose a patch-based data augmentation to model inter-scene correlation information at scene and instance levels simultaneously, facilitating information transfer between labeled scenes and unlabeled scene.
  • Extensive experiments demonstrate that our semi-supervised 3D segmentation methods achieves state-of-the-art performance, and comprehensive ablations studies are conducted to demonstrate its effectiveness.
@@
Figure 2: Overview of the proposed CoScene training pipeline. The teacher network generates the pseudo-labels $y_{i}^{u}$ from the unlabeled scene $x_{i}^{u}$, followed by an erasure operation to filter out low-confidence points, ensuring training consistency. For the labeled scene $x_{i}^{l}$ and the generated erased scene $\hat{x}_{i}^{u}$, we employ a patch-based data augmentation to obtain augmented scenes $x_{i}^{lpf}$ and $\hat{x}_{i}^{upf}$, separately. Subsequently, both scenes are fed into the student network for computing the loss of $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$.
Related Work
LiDAR Semantic Segmentation
LiDAR semantic segmentation aims to recognize the class label for each point of an input point cloud scene. Through the development of various deep neural networks, numerous approaches for LiDAR semantic segmentation have been explored over the past few years, which can be broadly grouped into three categories, i.e., voxel-based (Zhu et al. 2021), 2D projection-based (Kong et al. 2023a; Zhao, Bai, and Huang 2021), and point-based methods (Hu et al. 2020; Qi et al. 2017). Prior fully supervised works have made significant progress and achieved appealing performance, while the results deeply depend on large-scale precisely annotated datasets (Behley et al. 2019; Fong et al. 2022). To reduce annotations of 3D data, some methods with less supervision have been explored recently, e.g., weakly-supervised (Pan et al. 2024; Liu et al. 2022; Unal, Dai, and Van Gool 2022; Zhang et al. 2021; Chen et al. 2023; Tang et al. 2024) and semi-supervised learning (Li, Shum, and Breckon 2023; Reichardt, Ebert, and Wasenmüller 2023; Zou et al. 2018; Chen et al. 2021). LaserMix (Kong et al. 2023b), as the pioneer in semi-supervised LiDAR segmentation leverages the simple laser beam mix strategy to get solid results. Follow-up work fuses LiDAR-Camera data to mine deeper information (Chen et al. 2024). For semi-supervised approaches, the overall architecture often builds upon the mean-teacher paradigm (Tarvainen and Valpola 2017) and leverages the concept of pseudo label (Arazo et al. 2020). However, many points without pseudo labels contribute equally to pseudo-label points in the forward propagation, but are directly ignored when performing backward propagation, which brings intra-scene inconsistency issues. Hence, our CoScene leverages the point erasure strategy to mitigate this inconsistency problem, creating a complement to the pseudo-label mechanism to enable effective information transfer.
3D Point Cloud Data Augmentation
@@
Point cloud data augmentation has been explored across various 3D vision tasks (Zhu, Fan, and Weng 2023; Shin et al. 2023; Wang et al. 2024; Fang et al. 2021). Enlightened by the success of Mixup (Zhang et al. 2017) in image classification, several works explore the techniques for mixing selected point cloud samples. PointMixup (Chen et al. 2020) leverages interpolation to generate mixed samples for training. To preserve structural information, RSMix (Lee et al. 2021) provides a neighboring function, and PointCutMix (Zhang et al. 2022) establishes the correspondence relationship for sample fusion. Notably, these works aim at local object-level augmentation. As a result, several works also study the mixing strategies for point cloud scene augmentation. Mix3D (Nekrasov et al. 2021) concatenates points of two scenes to create a new training sample. PloarMix (Xiao et al. 2022) swaps points within the fan-shaped area between two scenes. CoSMix (Saltori et al. 2023) aims to mitigate the domain shift by mixing semantic point groups. These works only support similarity modeling between two scenes, but fail to construct correlation among multiple scenes. By contrast, our approach has the capability to model inter-scene information from multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enhance scenes. Therefore, our approach can effectively guide the transfer of annotation information from labeled scenes to unlabeled scenes, facilitating semi-supervised learning.
Method
In this section, we first state the definition of semi-supervised 3D semantic segmentation and introduce the mainstream teacher-student framework. Then, we present the intra-scene consistency and inter-scene correlation of CoScene, along with its core component, i.e., the point erasure strategy and patch-based data augmentation. Finally, we outline the overall training objective.
Preliminary
@@
Problem Formulation. Before we describe our model and training process, we first provide the definition of semi-supervised point cloud segmentation. In detail, we have a small labeled training set $D_{l}=\left\{x_{i}^{l},y_{i}^{l}\right\}_{i=1}^{N_{l}}$ and a large unlabeled training set $D_{u}=\left\{x_{i}^{u}\right\}_{i=1}^{N_{u}}$, where $x_{i}\in\mathbb{R}^{m\times{(3+r)}}$ is a point cloud scene of $m$ points with three-dimensional Cartesian coordinates relative to the sensor and additional $r$-dimensional feature (e.g., intensity, color, or reflectivity), $y_{i}$ represents its corresponding point-level segmentation labels, $N_{l}$ and $N_{u}$ are the numbers of labeled and unlabeled point cloud scenes, respectively, and $N_{u}\gg N_{l}$ in ideal cases. The goal of SSL-based point cloud segmentation is to learn a segmentation model using the limited labeled data and large-scale unlabeled data.
@@
Teacher-Student Paradigm. Following current mainstream research (Li, Shum, and Breckon 2023), our learning framework also builds upon the teacher-student paradigm, which includes two 3D segmentation networks with the same architecture. Here, we adopt Cylinder3D (Zhu et al. 2021) as the segmentation backbone $\Phi(\cdot)$ to enable a fair comparison with other methods. In detail, the teacher model generates pseudo-labels for unlabeled scenes and undergoes a gradual update through the exponential moving average (EMA) (Tarvainen and Valpola 2017) of the student weights:
$\theta_{\mathbb{t}}^{k+1}=\alpha\cdot\theta_{\mathbb{t}}^{k}+(1-\alpha)\cdot\theta_{\mathbb{s}}^{k},$ (1)
@@
where $\alpha$ is the EMA decay rate, which is set to 0.99 by default, $\theta_{\mathbb{t}}$ and $\theta_{\mathbb{s}}$ represent the weight parameters of the teacher and student models, respectively, and $k$ denotes the training step.
Point Erasure Strategy
@@
Based on the teacher-student mutual learning framework (Sohn et al. 2020), the overall architecture of CoScene is depicted in Figure 2. Specifically, CoScene consists of two primary branches: one dedicated to labeled scenes and the other to unlabeled scenes. How to effectively leverage a large number of unlabeled scenes is crucial for semi-supervised learning. For the branch of unlabeled scenes, we use a teacher network to generate pseudo-labeled scenes $(x_{i}^{u},y_{i}^{u})$ based on a predefined confidence threshold $\tau_{s}$ (Kong et al. 2023b). As mentioned previously, when using the pseudo-labeled scene for training, points without pseudo labels are involved in the forward propagation but are ignored in the backward propagation due to the lack of corresponding supervision. This leads to intra-scene inconsistency, which limits the effectiveness of semantic information transfer. To address this issue, we propose erasing points without pseudo labels, resulting in an erased scene $(\hat{x}_{i}^{u},\hat{y}_{i}^{u})$ that contains fully annotated information, similar to fully supervised scenes. Formally, intra-scene inconsistency addressed by the point erasure strategy can be defined as:
@@
where $\hat{x}_{i}^{u}$ is the erased scene with corresponding class labels $\hat{y}_{i}^{u}$, $\Phi_{s}(\cdot)$ represents the segmentation network that predicts the confidence score for each point in the scene $x_{i}^{u}$, and $\tau_{s}$ is a predefined confidence threshold, equivalent to the pseudo-label mining threshold, requiring no additional adjustment. Since the proposed point erasure strategy is independent of the segmentation network architecture, it can be utilized as a plug-and-play tool for semi-supervised 3D semantic segmentation methods based on the pseudo-label mechanism (see Figure 5 (a)).
Figure 3: Illustration of the pool initialization process. For each labeled point cloud scene, splitting it into multiple patches from the BEV perspective and assigning different instance categories to each patch enable the extraction of instances belonging to different classes. Subsequently, these patches and instances are stored individually in the patch pool and instance pool.
Patch-based Data Augmentation
Based on the labeled and erased scenes, we propose a patch-based data augmentation to enhance the transfer of annotation information, guided by inter-scene correlation. The key idea is to construct patch pools to store scene-level and instance-level information, and then combine patches from multiple scenes to generate diverse training samples.
Pool Initialization. We first use the labeled training set $D_{l}$ to create two patch pools, which store local-scene and single-instance information, respectively, as shown in Figure 3. Given the $i$-th point cloud scene $x_{i}^{l}$, we patchify $x_{i}^{l}$ into a sequence of $T$ regular, non-overlapping point patches $\{p_{ij}^{l}\}_{j=1}^{T}$, each of size $(n\times n)$, in the bird’s eye view. This process can be expressed as:

### v2

# Exploring Scene Affinity for Semi-Supervised LiDAR Semantic Segmentation

@@ Abstract
This paper explores scene affinity (AIScene), namely intra-scene consistency and inter-scene correlation, for semi-supervised LiDAR semantic segmentation in driving scenes. Adopting teacher-student training, AIScene employs a teacher network to generate pseudo-labeled scenes from unlabeled data, which then supervise the student network’s learning. Unlike most methods that include all points in pseudo-labeled scenes for forward propagation but only pseudo-labeled points for backpropagation, AIScene removes points without pseudo-labels, ensuring consistency in both forward and backward propagation within the scene. This simple point erasure strategy effectively prevents unsupervised, semantically ambiguous points (excluded in backpropagation) from affecting the learning of pseudo-labeled points. Moreover, AIScene incorporates patch-based data augmentation, mixing multiple scenes at both scene and instance levels. Compared to existing augmentation techniques that typically perform scene-level mixing between two scenes, our method enhances the semantic diversity of labeled (or pseudo-labeled) scenes, thereby improving the semi-supervised performance of segmentation models. Experiments show that AIScene outperforms previous methods on two popular benchmarks across four settings, achieving notable improvements of 1.9% and 5.3% in the most challenging 1% labeled data.
1 Introduction
Semantic segmentation of LiDAR point clouds is essential for autonomous driving, allowing vehicles to perceive their 3D surroundings. Most state-of-the-art models for driving-scene LiDAR segmentation are dependent on high-quality, point-annotated datasets [12, 55, 35]. With numerous objects in driving scenes, densely annotating points is extremely labor-intensive. Consequently, semi-supervised LiDAR semantic segmentation has attracted growing attention, where models are trained on a limited amount of labeled data supplemented by a large set of unlabeled data.
Figure 1: Top: Pseudo-labeled scenes generated from unlabeled scenes typically contain points with pseudo labels (colored) and without them (gray). Previous methods include all points in forward propagation but only pseudo-labeled points in backward propagation, leading to intra-scene inconsistency. This inconsistency may allow unsupervised, semantically ambiguous points to affect the learning of pseudo-labeled points (see Fig. 7). Bottom: Existing data augmentation techniques suitable for semi-supervised segmentation usually mix two scenes through scene-level operations like concatenation or swapping. Their resulting scenes may be constrained in terms of semantic diversity, potentially lacking objects like bicyclists and buses, which are important in driving scenes.
The common solution for learning from unlabeled data is to apply teacher-student training, where a teacher network generates pseudo-labeled scenes from unlabeled data, which are then used to supervise the learning of a student network. The student network’s objective is to minimize segmentation loss while the teacher network’s weights are updated via exponential moving average [38]. To further improve model learning with unlabeled data, existing studies implement designs such as utilizing language-vision models [22], adopting density-guided contrastive learning [21], or trying feature fusion between LiDAR and image data [5]. Despite the progress, two main issues remain to be addressed:
(1) Given the teacher network’s confident prediction, pseudo-labeled scenes typically contain points with and without pseudo-labels. During training, most studies include all points in pseudo-labeled scenes in the student network’s forward propagation, but only pseudo-labeled points are involved in backward propagation. Points excluded from backpropagation are unsupervised and, therefore, semantically ambiguous. However, these points still contribute to the forward feature computation for pseudo-labeled points, which may interfere with the model’s ability to learn from them effectively. In short, the inconsistency between forward and backward propagation within the scene causes ambiguous points to obstruct the model’s learning of pseudo-labeled points. Both point-based [12] and voxel-based [7, 55] methods suffer from this issue in feature computation, as illustrated at the top of Fig. 1.
(2) Data augmentation is an effective practice in semi-supervised LiDAR semantic segmentation [15, 13]. The basic idea is to mix raw labeled scenes and pseudo-labeled scenes to enrich scene semantics. Although impressive advancements have been made [32, 50], they usually select two scenes (one labeled scene and one pseudo-label scene) and then perform scene-level mix through operations like concatenation [26] and swapping [15], which limits the semantic diversity of mixed scenes. As shown at the bottom of Fig. 1, multiple scenes, whether labeled or pseudo-labeled, often share similar spatial layouts. Moreover, in driving scenes, objects such as cars and pedestrians are the primary focus of an autonomous vehicle’s attention. This naturally raises the question: can we improve the semantic diversity of scenes by mixing multiple scenes at the instance level?
To address inconsistency and implement the multi-scene mixing idea, we explore Scene AffInity for semi-supervised LiDAR semantic segmentation, named AIScene. The term scene affinity refers to intra-scene consistency and inter-scene correlation. Specifically, AIScene, built on the teacher-student training pipeline, employs a simple yet helpful point erasure strategy, removing points without pseudo-labels in forward propagation. This ensures intra-scene consistency in both forward and backward propagation, preventing semantically ambiguous points from disrupting learning from pseudo-labeled points. Building on this, AIScene incorporates patch-based data augmentation that mixes multiple scenes at both scene and instance levels. By maintaining two pools that store scene and instance patches from various scenes, it selects patches for mixing based on inter-scene correlation. This novel augmentation method not only enhances the semantic diversity of raw labeled scenes but also complements pseudo-labeled scenes where some points have been erased.
AIScene has been extensively evaluated through experiments and demonstrated superiority on two popular benchmarks for semi-supervised LiDAR semantic segmentation. It exceeds state-of-the-art methods by 1.3% on the SemanticKITTI [2] and 2.4% on the nuScenes [10] in terms of average performance across four settings (1%, 10%, 20%, and 50% labeled data). Notably, in the most challenging setting, with only 1% labeled data, AIScene achieves remarkable improvements of 1.9% and 5.3% on the two datasets. Additionally, our ablation studies confirm AIScene’s potential to reduce the requirements for labeled data across different LiDAR semantic segmentation networks (see Tab. 1 and  Tab. 2), with each component of AIScene able to function independently to facilitate semi-supervised LiDAR semantic segmentation (see Fig. 6).
To summarize, our contributions are as follows:
@@ 1 Introduction
  • We propose a point erasure strategy to ensure intra-scene consistency and facilitate network learning from pseudo-labeled scenes. This strategy can serve as a plug-and-play tool for semi-supervised LiDAR semantic segmentation using the pseudo-label mechanism.
  • We design a patch-based data augmentation to mix multiple scenes at both scene and instance levels, guided by inter-scene correlation. This method improves the semantics diversity of labeled scenes and complements pseudo-labeled scenes.
  • Extensive experiments demonstrate that AIScene achieves state-of-the-art results on popular benchmarks for semi-supervised LiDAR semantic segmentation. Ablation studies show that the novel erasure strategy and augmentation method can function independently.
@@
Figure 2: Overview of the proposed AIScene training pipeline. The teacher network generates the pseudo-labels $y_{i}^{u}$ from the unlabeled scene $x_{i}^{u}$, followed by an erasure operation to filter out low-confidence points, ensuring training consistency. For the labeled scene $x_{i}^{l}$ and the generated erased scene $\hat{x}_{i}^{u}$, we employ patch-based data augmentation to obtain augmented scenes $x_{i}^{lpf}$ and $\hat{x}_{i}^{upf}$, respectively. Subsequently, both scenes are fed into the student network to compute the loss of $\mathcal{L}_{l}$ and $\mathcal{L}_{u}$.
2 Related Work
Fully-supervised LiDAR Semantic Segmentation: Numerous approaches for LiDAR semantic segmentation have been explored, which can be broadly grouped into three categories: voxel-based, 2D projection-based, and point-based methods. Voxel-based approaches [6, 7, 36, 55, 19] involve voxelizing sparse point clouds to generate regular voxel grids, allowing the application of standard 3D convolution neural networks for embedding learning. Conversely, 2D projection-based methods [14, 25, 43, 46, 51, 53] project input point clouds into a 2D dense space, use 2D convolutional filters [30] for processing, and then map predictions back to the original 3D space. Finally, point-based methods [12, 28, 18, 44] operate directly on raw point clouds, enabling the extraction of point-wise semantic features.
Semi-supervised LiDAR Semantic Segmentation: To reduce the need for 3D dense annotations, recent research has explored methods that require less supervision [17, 24, 39, 52, 3, 37]. Semi-supervised learning [23, 29, 56, 4, 21, 16] aims to utilize a large number of unlabeled data to alleviate the annotation burden. LaserMix [15], a pioneer work in semi-supervised LiDAR segmentation, leverages a simple laser beam mix strategy to achieve solid results. Subsequent research fuses LiDAR and camera data to extract deeper semantic information [5, 40]. DDSemi [21], focusing on point cloud density, builds on contrastive learning to enhance feature discriminability. The recently proposed LASS3D [22] leverages large language-vision models to support semi-supervised segmentation. Existing semi-supervised methods often build upon the mean-teacher paradigm [38] and utilize pseudo labels [1]. However, many points without pseudo-labels contribute equally to pseudo-labeled points in forward propagation but are ignored during backward propagation, leading to intra-scene inconsistency. To address this, our AIScene leverages a point erasure strategy to mitigate this inconsistency, complementing the pseudo-label mechanism.
@@ 2 Related Work
3D Point Cloud Data Augmentation: Point cloud data augmentation has been explored across various 3D vision tasks [54, 33, 42, 9]. Inspired by the success of Mixup [49] in image classification, several methods have investigated techniques for mixing selected point cloud samples. To preserve structural information, RSMix [20] introduces a neighboring function, and PointCutMix [50] establishes correspondences for sample fusion. Notably, these methods focus on local object-level augmentation. In response, several works have explored mixing strategies for point cloud scene augmentation. PloarMix [45] swaps points within the fan-shaped areas between two scenes, while CoSMix [31] aims to mitigate domain shift by mixing semantic point groups. Unlike these methods, our method is capable of modeling inter-scene information across multiple scenes to create more diversified data, while also extracting instance-level semantic information for individual classes to further enrich scenes.
3 Method
3.1 Preliminary
@@
Problem Formulation. Before we describe our model and training process, we first provide the definition of semi-supervised point cloud segmentation. In detail, we have a small labeled training set $D_{l}=\left\{x_{i}^{l},y_{i}^{l}\right\}_{i=1}^{N_{l}}$ and a large unlabeled training set $D_{u}=\left\{x_{i}^{u}\right\}_{i=1}^{N_{u}}$, where $x_{i}\in\mathbb{R}^{m\times{(3+r)}}$ is a point cloud scene of $m$ points with three-dimensional Cartesian coordinates relative to the sensor and additional $r$-dimensional feature (e.g., intensity, color, or reflectivity), $y_{i}$ represents its corresponding point-level segmentation labels, $N_{l}$ and $N_{u}$ are the numbers of labeled and unlabeled point cloud scenes, respectively, and $N_{u}\gg N_{l}$ in ideal cases. The goal of SSL-based point cloud segmentation is to learn a segmentation model using limited labeled data and large-scale unlabeled data.
@@
Teacher-Student Paradigm. Following current mainstream research [15, 21], our framework also builds upon the teacher-student paradigm, which includes two segmentation networks with the same architecture. Our AIScene can be instantiated with any 3D segmentation network [55, 7, 19], denoted as $\Phi(\cdot)$. In detail, the teacher model generates pseudo-labels for unlabeled scenes and undergoes a gradual update through the exponential moving average (EMA) [38] of the student weights:
$\theta_{\mathbbm{t}}^{k+1}=\alpha\cdot\theta_{\mathbbm{t}}^{k}+(1-\alpha)\cdot\theta_{\mathbbm{s}}^{k},$ (1)
@@
where $\alpha$ is the EMA decay rate, which is set to 0.99 by default, $\theta_{\mathbbm{t}}$ and $\theta_{\mathbbm{s}}$ represent the weight parameters of the teacher and student models, respectively, and $k$ denotes the training step.
3.2 Point Erasure Strategy
@@
Based on the teacher-student mutual learning framework [34], the overall architecture of AIScene is depicted in Fig. 2. Specifically, AIScene consists of two primary branches: one dedicated to labeled scenes and the other to unlabeled scenes. How to effectively leverage a large number of unlabeled scenes is crucial for semi-supervised learning. Hence, for the branch of unlabeled scenes, we use a teacher network to generate pseudo-labeled scenes $(x_{i}^{u},y_{i}^{u})$ based on a predefined confidence threshold $\tau_{s}$ [15]. As mentioned previously, when using the pseudo-labeled scene for training, points without pseudo-labels are involved in the forward propagation but are ignored in the backward propagation due to the lack of corresponding supervision. This leads to intra-scene inconsistency, which causes semantic ambiguity of points. To address this issue, we propose erasing points without pseudo labels, resulting in an erased scene $(\hat{x}_{i}^{u},\hat{y}_{i}^{u})$ that contains fully annotated information, similar to fully supervised scenes. Formally, intra-scene inconsistency addressed by the point erasure strategy can be defined as:
@@
where $\hat{x}_{i}^{u}$ is the erased scene with corresponding class labels $\hat{y}_{i}^{u}$, $\Phi_{s}(\cdot)$ represents the segmentation network that predicts the confidence score for each point in the scene $x_{i}^{u}$, and $\tau_{s}$ is a predefined confidence threshold, equivalent to the pseudo-label mining threshold, requiring no additional adjustment. Since the proposed point erasure strategy is independent of the segmentation network architecture, it can be utilized as a plug-and-play tool for semi-supervised 3D semantic segmentation methods based on the pseudo-label mechanism, as depicted in Fig. 6 (a).
Figure 3: Illustration of the pool initialization process. For each labeled point cloud scene, splitting it into multiple patches from the BEV perspective and assigning different instance categories to each patch enables the extraction of instances belonging to different classes. Subsequently, these patches and instances are stored individually in the patch pool and instance pool.
3.3 Patch-based Data Augmentation
Based on the labeled and erased scenes, we propose a patch-based data augmentation to enhance the semantic diversity of annotation information, guided by inter-scene correlation. The key idea is to construct patch pools to store scene-level and instance-level information, and then combine patches from multiple scenes to generate diverse training samples.
Pool Initialization. We first use the labeled training set $D_{l}$ to create two patch pools, which store scene-level and instance-level information, respectively, as shown in Fig. 3. Given the $i$-th point cloud scene $x_{i}^{l}$, we patchify $x_{i}^{l}$ into a sequence of $T$ regular, non-overlapping point patches $\{p_{ij}^{l}\}_{j=1}^{T}$, each of size $(n\times n)$, in the bird’s eye view. This process can be expressed as: